OpenStudy (anonymous):

Write a polynomial that represents the product of the integer before x and the next consecutive odd integer after x

OpenStudy (anonymous):

x(x - 1) x(x + 1) (x - 1)(x + 1) (x + 1)(x + 1)

OpenStudy (mathstudent55):

x is an integer. The integer before x is 1 less than x. How would you represent 1 less than x in terms of x?

OpenStudy (anonymous):

x(x - 1)

OpenStudy (mathstudent55):

I'm not asking for an answer form the choices. We are only looking for one of the numbers now. What is 1 less than x?

OpenStudy (anonymous):

0

OpenStudy (mathstudent55):

No. Let me try with numbers. A number is 8. What is 1 less than 8?

OpenStudy (anonymous):

7

OpenStudy (mathstudent55):

Correct. What operation did you do to 8 to get the correct answer of 7?

OpenStudy (anonymous):

subtraction

OpenStudy (mathstudent55):

Correct. 1 less than 8 means 8 - 1 = 7 1 less than 20 means 20 - 1 = 19 1 less than 100 means 100 - 1 = 99 You follow all those examples?

OpenStudy (anonymous):

yep

OpenStudy (mathstudent55):

Now thinking the same way, that 1 less than a number means subtract one form the number, what is 1 less than x? It is x - 1 We can't simplify x - 1 any more because x and 1 are not like terms. x - 1 is simply x - 1 and it means whatever number you allow x to be, subtract 1 from it.

OpenStudy (mathstudent55):

Now let's look at the problem again.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Write a polynomial that represents the product of the integer before x and the next consecutive odd integer after x

OpenStudy (mathstudent55):

The integer before x is 1 less than x. That means the integer before x is x - 1. We already know that.

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

The next consecutive integer after x is 1 more than x. Just like 1 less than x is x - 1, 1 more than x is x + 1. To show the product of one less than x and one more than x, you do (x - 1)(x + 1)

OpenStudy (anonymous):

so its(x - 1)(x + 1)?

OpenStudy (mathstudent55):

Yes.